OpenStudy (anonymous):

Solve the logarithm In x = 2 Solve the logarithm In x = 2 @Mathematics

OpenStudy (anonymous):

7.389

OpenStudy (anonymous):

ln is the same as log base e

OpenStudy (anonymous):

or e^2

OpenStudy (anonymous):

0.6931

OpenStudy (anonymous):

no, .6391 yields a negative answer

OpenStudy (anonymous):

with base 10,it is,0.301

OpenStudy (anonymous):

What you got was ln(2) which is not right

OpenStudy (turingtest):

do you get what to do in these problems? they are very simple: raise e to the power of both sides e^(lnx)=e^2 x=e^2

jhonyy9 (jhonyy9):

x=e squared

OpenStudy (anonymous):

The value of log2 depends on the base of the logarithm. log without a subscript usually measn base 10, but is sometimes used for base e. base 10: log2 = 0.3010299956 ≈ 0.30103 base e: log2 = 0.69314718

OpenStudy (anonymous):

No, we need to solve for x in ln(x)=2

OpenStudy (anonymous):

thn,it is e2

OpenStudy (anonymous):

Thank you, can someone show the work? cause i dont understand how to exactly do this kind of lagarithm

jhonyy9 (jhonyy9):

no lg signe logarithm on base 10

OpenStudy (anonymous):

Check tourist Test's post, he explained it

OpenStudy (anonymous):

Test showed the work a bit up

jhonyy9 (jhonyy9):

check wikipedia

OpenStudy (turingtest):

\[\ln x = 2 \]\[e^{\ln x}=e^2\]\[x=e^2\]

jhonyy9 (jhonyy9):

on wikipedia

OpenStudy (turingtest):

if it were base 5 for example:\[\log_5x=2\]\[5^{\log_5x}=5^2\]\[x=5^2=25\]just watch the bases

OpenStudy (anonymous):

Thank you everyone, and thank you soo much turingtest for the explanation :D

OpenStudy (turingtest):

no prob :)

OpenStudy (anonymous):

You are the best, and helped me really understand when you helped me with all my problems lol ^^

OpenStudy (turingtest):

I try!

OpenStudy (anonymous):

Hehe thank you again :)